### v2

Abstract
Integration of unmanned aerial vehicles (UAVs) for surveillance or monitoring applications into fifth generation (5G) New Radio (NR) cellular networks is an intriguing problem that has recently tackled a lot of interest in both academia and industry. For an efficient spectrum usage, we consider a recently-proposed sky-ground nonorthogonal multiple access (NOMA) scheme, where a cellular-connected UAV acting as aerial user (AU) and a static terrestrial user (TU) are paired to simultaneously transmit their uplink signals to a base station (BS) in the same time-frequency resource blocks. In such a case, due to the highly dynamic nature of the UAV, the signal transmitted by the AU experiences both time dispersion due to multipath propagation effects and frequency dispersion caused by Doppler shifts. On the other hand, for a static ground network, frequency dispersion of the signal transmitted by the TU is negligible and only multipath effects have to be taken into account. To decode the superposed signals at the BS through successive interference cancellation, accurate estimates of both the AU and TU channels are needed. In this paper, we propose channel estimation procedures that suitably exploit the different circular/noncircular modulation formats (modulation diversity) and the different almost-cyclostationarity features (Doppler diversity) of the AU and TU by means of widely-linear time-varying processing. Our estimation approach is semi-blind since Doppler shifts and time delays of the AU are estimated based on the received data only, whereas the remaining relevant parameters of the AU and TU channels are acquired relying also on the available training symbols. Monte Carlo numerical results demonstrate that the proposed channel estimation algorithms can satisfactorily acquire all the relevant parameters in different operative conditions.
{IEEEkeywords}
Almost-cyclostationarity, channel estimation, circular and noncircular modulations, double-selective channels, Doppler diversity, 5G New Radio (NR), modulation diversity, non-orthogonal multiple access (NOMA), unmanned aerial vehicles (UAVs), uplink.
1 Introduction
[…]
With the development of unmanned aerial vehicles (UAVs), computer vision, and sensor technology, UAV systems have been increasingly employing in civilian and commercial applications, such as surveillance and monitoring, due to their ability to quickly cover large and difficult-to-reach areas. The recently-introduced concept of cellular-connected UAVs [1, 2] represents a viable strategy to integrate UAVs in fifth generation (5G) New Radio (NR) networks, by allowing them to transmit data to the cellular network as aerial users (AUs). 3GPP Release 18 is introducing 5G NR support to enhance the safe use of UAVs for commercial and leisure applications. Nevertheless, the integration of UAVs into cellular networks introduces two interelated basic challenges [3, 4]. First, existing 5G cellular infrastructures are designed to serve terrestrial users (TUs) and, thus, the antennas of the base stations (BSs) look downward: such a tilting might cause AUs to be served by the side lobes of the BSs and, consequently, they suffer from a reduction in antenna gain. Second, due to their mobility capabilities, AUs have more versatile movements than TUs, thereby introducing Doppler (frequency) shifts, which result in carrier frequency offset, inter-carrier interference, and reduced channel coherence time.
[…]
The feasibility of using existing cellular networks to support UAVs in the low-altitude airspace has been demonstrated in [5, 6, 7]. Specifically, it has been shown that the favorable line-of-sight (LoS) propagation conditions for UAVs flying in the sky can compensate for the reduced gain of antenna side lobes of the BSs, provided that UAVs fly below $120-200$ m. A first consequence of flying at lower altitudes is that the AU-to-BS channel exhibits multipath components (MPCs) consisting of an LoS link with high probability and a cluster of reflected, delayed paths [8]. Severe Doppler shifts are caused by the high carrier frequency and high velocity of the UAV, and they are influenced by the angular distribution of the scattered components. Indeed, different MPCs may have largely different Doppler shifts.
To support AUs in cellular networks, there are two possible multiple access schemes. In the case of orthogonal multiple access (OMA) [9, 10, 11, 12, 13, 14, 15, 16, 17, 18], AUs and TUs are served using orthogonal resource blocks (RBs). On the other hand, non-orthogonal multiple access (NOMA) allows AUs and TUs to simultaneously share the same spectrum through power-domain or code-domain multiplexing: in this case, successive interference cancellation (SIC) is used at the BS to separate the signals. In many cases, NOMA ensures superior spectral efficiency with respect to OMA [19, 20, 21, 22, 23, 24].
1.1 NOMA for cellular-connected UAVs
Outage probabilities in downlink and uplink NOMA transmissions have been evaluated in [25] based on instantaneous distinct signal power for devices including UAVs. In [26], an aerial-ground NOMA scheme that pairs the AU and TU for data and control links has been investigated, by exploiting the asymmetric features of the channels and rate demands of the AU and TU in the downlink communication. The authors in [27] have proposed a cooperative downlink NOMA scheme to cancel co-channel interference via SIC in ideal two-cell networks. A cooperative NOMA scheme that exploits existing backhaul links among BSs for SIC operations has been proposed in [28]. An uplink NOMA has been proposed in [29] to serve AU and TU, where AU trajectory and its cell-association order are jointly optimized, whereas the use of uplink NOMA for AUs and TUs has been studied via stochastic geometry, where the mobility of the AUs is taken into account. The probability that the achievable data rate of both the AU and TU exceeds the respective target rates is calculated in [30], where the minimum height that the AU needs to fly, at each transmission point along the given trajectory, is also numerically determined in order to satisfy a certain quality of service constraint. Finally, the authors in [31] have derived the optimal precoding that maximizes the uplink sum-rate of an AU and multiple TUs.
A common underlying approach of all the aforementioned works [25, 26, 27, 28, 29, 30, 31] is that Doppler shifts due to the UAV motions are a cause of performance degradation rather than a source of diversity. Indeed, it is tacitly assumed that large Doppler shifts have been previously compensated for or, as in [30], that the AU transmits its data to the BS by maintaining a fixed position over selected points along a given trajectory (i.e., transmission during hovering flight). However, compensation of Doppler shifts is a challenging task in multipath channels, except for the case when the MPCs of the AU-to-BS channel have a very similar Doppler shift, which is a situation that may not be fulfilled in the low-altitude airspace [8]. Moreover, allowing the UAV to transmit only in the hovering-flight state along its trajectory entails a waste of both communication and power resources. Another shortcoming of the aforementioned papers [25, 26, 27, 28, 29, 30, 31] is that channel estimation issues have not been accounted at all. In a NOMA setting, channel acquisition of both the AU and TU channels is crucial to obtain accurate SIC at the BS.
The works [32, 33] deserve a separate discussion. Orthogonal time frequency space (OTFS) modulation has been used in [32] for the implementation of NOMA among users with different mobility profiles: specifically, a high-mobility user is allowed to share the spectrum with multiple low-mobility NOMA users. Although the specific case of an AU is not explicitly considered in [32, 33], their framework may be adapted to the scenario of cellular-connected UAVs, by regarding the high-mobility user as an UAV. However, besides entailing more complex signal processing techniques at both the transmitter and receiver, compared to standard orthogonal frequency-division multiplexing (OFDM) modems, zero fractional delays and zero fractional Doppler shifts are assumed in the OTFS-NOMA scheme. Last but not least, Doppler shifts are still seen as a source of performance degradation to be counteracted and perfect channel knowledge is assumed in [32, 33].
Channel estimation specifically targeted at cellular-connected UAVs remains a critical issue, and effective methods operating over shared aerial and terrestrial channels are an open research issue.
1.2 Contribution and organization
We consider a NOMA scheme, which is referred to as sky-ground (SG) NOMA, where a flying AU is paired with a static (i.e., fixed) TU. The AU and TU have different propagation conditions and their received power can be rather different at the BS, regardless of their transmit power: indeed, the air-to-cellular (A2C) channel between the AU and the elevated terrestrial BS is both time- and frequency-selective (i.e., doubly-selective) with significant Doppler shifts, [^1] The air-to-cellular channel is different from the air-to-ground one due to the non-negligible height of terrestrial BSs. whereas the TU-to-BS channel is predominately frequency-selective with negligible Doppler effects. We propose to exploit the different mobile profiles of the AU and TU rather than to counteract the Doppler effects induced by the mobility of the UAV. In particular, the AU is allowed to continuously transmit along all its trajectory.
To counteract the frequency-selectivity of both channels, the AU and TU resort to the OFDM transmission format. To improve the user separation capability of the BS, we propose to exploit a further source of diversity between the two users in the modulation domain: in both training (or pilot) and data phases, the TU transmits circular symbols, whereas the AU employs noncircular modulation.
Acquisition of channel state information (CSI) in uplink NOMA is a challenging task, since channel estimation in current wireless data communications standards is mainly based on training sequences sent from the users to the BS that are orthogonal in either time or frequency domain. Orthogonal RBs for training adversely affect spectral efficiency and system performance, especially in the uplink. We consider a training scheme where the training data of one user is contaminated by either the pilots or the information-bearing data of the other user. In this challenging scenario, we propose a two-stage approach: first, the channel parameters of the AU are semi-blindly estimated by also exploiting the almost-cyclostationarity (ACS) properties [34] of the received signal that depend on the Doppler shifts of the AU-to-BS channel (Doppler diversity), as well as the different circular/noncircular features of the AU and TU (modulation diversity); subsequently, capitalizing on the CSI knowledge of the AU, the overall channel of the TU is acquired with a pilot-aided improved widely-linear (WL) estimator [35]. The performances of the proposed channel estimation algorithms are validated through extensive numerical examples.
The paper is organized as follows. The system model of the considered uplink SG-NOMA is described in Section 2. The proposed estimation strategies for the AU and TU are developed in Sections 3 and 4, respectively. Numerical results are reported in Section 5 in terms of channel estimation accuracy. Finally, the main results obtained in the paper are summarized in Section 6.
1.3 Basic notations
Upper- and lower-case bold letters denote matrices and vectors; the superscripts ***, T, H, and $-1$ denote the conjugate, the transpose, the Hermitian (conjugate transpose), and the inverse of a matrix; $\mathbb{C}$, $\mathbb{R}$ and $\mathbb{Z}$ are the fields of complex, real and integer numbers; $\mathbb{C}^{n}$ $[\mathbb{R}^{n}]$ denotes the vector-space of all $n$-column vectors with complex [real] coordinates; similarly, $\mathbb{C}^{n\times m}$ $[\mathbb{R}^{n\times m}]$ denotes the vector-space of all the $n\times m$ matrices with complex [real] elements; $\delta(\tau)$ is the Dirac delta; $\delta_{n}$ is the Kronecker delta, i.e., $\delta_{n}=1$ when $n=0$ and zero otherwise; $\jmath\triangleq\sqrt{-1}$ denotes the imaginary unit; $\text{max}(x,y)$ [$\text{min}(x,y)$] returns the maximum [minimum] between $x\in\mathbb{R}$ and $y\in\mathbb{R}$; $\lfloor x\rfloor$ rounds $x\in\mathbb{R}$ to the nearest integer less than or equal to $x$; the (linear) convolution operator is denoted with $\ast$; $\frac{\partial}{\partial x*}f$ is the complex derivative of the real-valued scalar function $f$ with respect to $x^{*}$ [36]; $\mathbf{0}_{n}$, ${\mathbf{O}}_{n\times m}$ and ${\mathbf{I}}_{n}$ denote the $n$-column zero vector, the $n\times m$ zero matrix and the $n\times n$ identity matrix; ${\mathbf{J}}_{n}$ is given by
${\mathbf{J}}_{n}\triangleq\begin{pmatrix}{\mathbf{O}}_{n\times n}&{\mathbf{I}}_{n}\\ {\mathbf{I}}_{n}&{\mathbf{O}}_{n\times n}\end{pmatrix}\in\mathbb{R}^{(2n)\times(2n)}\>;$ (1)
[…]
1.4 Special functions and matrices
The Dirichlet (or periodic sinc) function is defined as
${\rm D}_{n}(x)\triangleq\frac{\sin(\pi xn)}{\sin(\pi x)}\,e^{-j\pi(n-1)x}$ (2)
for each $x\in\mathbb{R}$. ${\mathbf{W}}_{n}\in\mathbb{C}^{n\times n}$ is the unitary symmetric $n$-point inverse discrete Fourier transform (IDFT) matrix, whose $(m+1,p+1)$-th entry is given by $\frac{1}{\sqrt{n}}\,e^{\jmath\frac{2\pi}{n}mp}$ for $m,p\in\{0,1,\ldots,n-1\}$, and its inverse ${\mathbf{W}}_{n}^{-1}={\mathbf{W}}_{n}^{\text{H}}$ is the $n$-point discrete Fourier transform (DFT) matrix; ${\mathbf{F}}\in\mathbb{R}^{P\times P}$ and ${\mathbf{B}}\triangleq{\mathbf{F}}^{\text{T}}\in\mathbb{R}^{P\times P}$ denote the forward shift and backward shift matrices, whose first column and the first row are given by $[0,1,0,\ldots,0]^{\text{T}}$ and $[0,1,0,\ldots,0]$, respectively.
2 System model and basic assumptions
The considered single-cell network scenario encompasses an elevated BS, mobile AUs, and static TUs. The BS is located within the cell at a fixed height $H_{\text{BS}}$, whereas the AU flies at an altitude $H_{\text{A}}$ above the ground. [^2] The proposed estimation algorithms can be straightforwardly modified to account for the case where the AU varies its height along the trajectory. We restrict ourselves to uplink communications since advanced signal detection algorithms of NOMA are more affordable at the BS [24]. A uniform linear array (ULA) is used at the BS with $J$ antennas. The AUs are represented by single-antenna UAVs that are employed for surveillance or monitoring a given area within a cell and transmit data to the corresponding BS over given A2C links. According to existing channel access schemes for terrestrial cellular networks, we assume that TUs transmit to the BS using the OMA scheme and, hence, there is no interference among the TUs in the cell.
In this paper, we consider an uplink SG-NOMA protocol, where single-antenna TUs share their RBs with the AUs. As in [30], we assume that each AU is randomly paired with a different static (i.e., fixed) TU (see Fig. 1). The two-user scenario represents the case of major practical interest for NOMA, since it results in acceptable interference levels [38].
Figure 1: Uplink sky-ground NOMA: the UAV and the terrestrial user transmit their own uplink signals to the BS in the same RBs.
2.1 Transmit signal model
Both users employ OFDM transmission to cope with the time-dispersive nature of the channel.
The independent symbol streams emitted by the AU and the TU ($n\in\mathbb{Z}$) are denoted with $\{s_{\text{A}}[n]\}$ and $\{s_{\text{T}}[n]\}$, respectively. [^3] Throughout the paper, the subscripts A and T refer to the AU (i.e, UAV) and the TU, respectively. We model $\{s_{\text{A}}[n]\}$ as a sequence of zero-mean unit-variance independent and identically distributed (i.i.d.) complex noncircular random variables [39], with second-order moment ${\mathbb{E}}(s_{\text{A}}^{2}[n])\neq 0$. On the other hand, $\{s_{\text{T}}[n]\}$ is modeled as a sequence of zero-mean unit-variance i.i.d. complex circular random variables, i.e., ${\mathbb{E}}(s_{\text{T}}^{2}[n])=0$. 5G NR has been natively designed for circular modulation schemes, such as phase shift keying (PSK) and quadrature amplitude modulation (QAM). Noncircular modulation schemes have been subsequently introduced for the uplink data channel (physical uplink shared channel- PUSCH) and control channel (physical uplink control channel - PUCCH) transmissions [40, 41, 42], such as $\pi/2$ binary PSK (BPSK). These waveforms, when combined with an appropriate spectrum shaping, enable low peak-to-average-power ratio transmissions without compromising the error rate performance, which is a desirable property for UAV communications. We will show herein that exploitation of the different second-order statistics (i.e, noncircular versus circular) of the AU and TU signals is instrumental into acquiring CSI in NOMA.
Both users share the same $M$ orthogonal subcarriers. Let $s_{\text{TX}}^{[m]}[\ell]\triangleq s_{\text{TX}}[\ell\,M+m]$ represent the symbol transmitted in the $\ell$th data block on the $m$th subcarrier, with $\text{TX}\in\{\text{A},\text{T}\}$, the vector ${\mathbf{s}}_{\text{TX}}[n]\triangleq(s_{\text{TX}}^{[0]}[n],s_{\text{TX}}^{[1]}[n],\ldots,s_{\text{TX}}^{[M-1]}[n])^{\text{T}}\in\mathbb{C}^{M}$ is subject to IDFT and cyclic prefix (CP) insertion, thus yielding ${\mathbf{u}}_{\text{TX}}[n]\triangleq(u_{\text{TX}}^{[0]}[n],u_{\text{TX}}^{[1]}[n],\ldots,u_{\text{TX}}^{[P-1]}[n])^{\text{T}}={\mathbf{I}}_{\text{cp}}\,{\mathbf{W}}_{M}\,{\mathbf{s}}_{\text{TX}}[n]$, where ${\mathbf{I}}_{\text{cp}}\in\mathbb{R}^{P\times M}$ models the insertion of a CP of length $L_{\text{cp}}$, with $P\triangleq M+L_{\text{cp}}$, and ${\mathbf{W}}_{M}$ is the $M$-point IDFT (see Subsection I-D). The precoded data vector ${\mathbf{u}}_{\text{TX}}[n]$ undergoes parallel-to-serial (P/S) conversion, and the resulting data sequence $\{u_{\text{TX}}[\ell]\}$ feeds a digital-to-analog converter (DAC), operating at rate $1/T_{\text{c}}=P/T_{\text{s}}$, where $u_{\text{TX}}[\ell\,P+q]=u_{\text{TX}}^{[q]}[\ell]$, for $q\in\{0,1.\ldots,P-1\}$, the OFDM symbol period is $T_{\text{s}}$, whereas $T_{\text{c}}$ is the sampling period. The baseband transmitted continuous-time signal at the output of the DAC is given by
$x_{\text{TX}}(t)=\sum_{\ell=-\infty}^{+\infty}\sum_{q=0}^{P-1}u_{\text{TX}}^{[q]}[\ell]\,\psi_{\text{DAC}}(t-q\,T_{\text{c}}-\ell\,T_{\text{s}})$ (3)
for $\text{TX}\in\{\text{A},\text{T}\}$, where $\psi_{\text{DAC}}(t)$ denotes the impulse response of the DAC.
2.2 Channel model for terrestrial user
The far-field channel between the TU and the BS is invariant over a coherence time $T_{\text{coh}}\triangleq N_{\text{coh}}\,T_{\text{s}}$, with $N_{\text{coh}}\gg 1$, wherein it can be modeled as a linear time invariant (LTI) system with impulse response:
${\mathbf{h}}_{\text{T}}(\tau)=\sum_{k=1}^{K_{\text{T}}}g_{\text{T},k}\,\delta(\tau-\tau_{\text{T},k})\,{\mathbf{a}}(\theta_{\text{T},k})$ (4)
where $K_{\text{T}}$ is the number of paths between the terrestrial user and the BS, whereas $g_{\text{T},k}$, $\tau_{\text{T},k}$, and $\theta_{\text{T},k}$ denote the complex gain, the time delay, and the angle of arrival (AoA) at the BS of the $k$th path. Taking the first receive antenna as reference, the vector ${\mathbf{a}}(\theta)\in\mathbb{C}^{J}$ in (4) denotes the response vector of the BS in the direction $\theta\in[-\pi/2,\pi/2]$ relative to the boresight of the ULA and it is given by
${\mathbf{a}}(\theta)\triangleq\left(1,e^{\jmath\frac{2\pi}{\lambda_{\text{carrier}}}d\sin(\theta)},\ldots,e^{\jmath\frac{2\pi}{\lambda_{\text{carrier}}}(J-1)d\sin(\theta)}\right)^{\text{T}}\in\mathbb{C}^{J}$ (5)
where $d$ is the distance between two adjacent antennas and $\lambda_{\text{carrier}}=c/f_{\text{carrier}}$ is the wavelength, with $c$ being the light speed. From now on, we will set $d=\lambda_{\text{carrier}}/2$ for simplicity.
Since each propagation path is approximately equal to the sum of independent micro-scatterers contributions, having same time delay and AoA, the channel gains $\{g_{\text{T},k}\}$ can thus be modeled as independent zero-mean complex Gaussian random variables (Rayleigh fading model), with variances $\sigma_{\text{T},k}^{2}\triangleq{\mathbb{E}}(|g_{\text{T},k}|^{2})$, for $k\in\{1,2,\ldots,K_{\text{T}}\}$. It can be adopted for $\sigma_{\text{T},k}^{2}$ the common propagation model (see, e.g., [43])
$\sigma_{\text{T},k}^{2}=\sigma_{\text{T,ref}}^{2}\,\mathcal{P}_{\text{T}}\,G_{\text{T}}\,G_{\text{BS,T}}\,\,\left(\frac{d_{\text{T,ref}}}{d_{\text{T},k}}\right)^{\eta_{\text{T}}}$ (6)
where $\sigma_{\text{T,ref}}^{2}$ is a unitless constant, $\mathcal{P}_{\text{T}}$ is the transmit power of the TU, $G_{\text{T}}$ and $G_{\text{BS,T}}$ represent the antenna gains of the TU and the BS, respectively, $d_{\text{T},k}$ denotes the propagation length along the $k$-th path, $d_{\text{T,ref}}$ is a reference distance for the antenna far field, and $\eta_{\text{T}}$ is the path-loss exponent.
2.3 Channel model for the aerial user
For A2C link, we assume a two-ray channel model, which includes the LoS path and a scattered (i.e., non-LoS) component [8] (see also Fig. 1). The non-LoS (NLoS) component is physically due to many rays reflected or scattered from closely-spaced terrestrial obstacles within the vicinity of the BS, which are characterized by a very narrow beamwidth [8] and appear grouped or “clustered" in delay for sufficiently high elevation angles [44]. Henceforth, such rays are expected to have a very similar Doppler shift, as they all come from similar directions (narrow beamwidth), and a very similar delay (high elevation angle). Assuming that the UAV is moving (relative to the BS) at constant radial speed $v$ within the observation interval $[0,T_{0})$, with $T_{0}\triangleq N_{0}\,T_{\text{s }}\leq T_{\text{coh}}$, under the customary assumption that the communication bandwidth $B\sim 1/T_{\text{c}}$ is much smaller than the carrier frequency $f_{\text{carrier}}$, the far-field low-pass equivalent linear time-varying (LTV) time-delay impulse response of the UAV-to-BS channel can be expressed as [^4] Throughout the paper, the subscripts L and N indicate parameters referring to the LoS and NLoS components of the AU, respectively. (see, e.g., [45])
${\mathbf{h}}_{\text{A}}(t,\tau)=g_{\text{A,L}}\,e^{\jmath\,2\pi f_{\text{A,L}}t}\,\delta(\tau-\tau_{\text{A,L}})\,{\mathbf{a}}(\theta_{\text{A,L}})\\ +g_{\text{A,N}}\,e^{\jmath\,2\pi f_{\text{A,N}}t}\,\delta(\tau-\tau_{\text{A,N}})\,{\mathbf{a}}(\theta_{\text{A,N}})$ (7)
where $g_{\text{A,L}}$ [$g_{\text{A,N}}$], $\theta_{\text{A,L}}$ [$\theta_{\text{A,N}}$], $\tau_{\text{A,L}}$ [$\tau_{\text{A,N}}$], $f_{\text{A,L}}=f_{\text{max}}\,\cos(\vartheta_{\text{A,L}})$ [$f_{\text{A,N}}=f_{\text{max}}\,\cos(\vartheta_{\text{A,N}})$] are the complex (i.e., amplitude and phase) path gain, the AoA at the BS, the time delay, and the Doppler shift of the LoS [NLoS] component, with $\vartheta_{\text{A,L}}\in[0,\pi]$ [$\vartheta_{\text{A,N}}\in[0,\pi]$] representing the angle of departure (AoD) of the LoS [NLoS] ray relative to the direction of motion of the UAV and $f_{\text{max}}\triangleq{f_{\text{carrier}}\,v}/{c}\ll B$. It is noteworthy that the LTV model (7) is justified by the fact that the UAV is allowed to communicate with the BS while it is flying and, hence, we do not constrain the UAV to transmit only in the hovering-flight state as, e.g., in [30].
Since the speed and physical location of the UAV would change much slower than the symbol period $T_{\text{s}}$, the UAV related parameters $g_{\text{A,L}}$, $g_{\text{A,N}}$, $\theta_{\text{A,L}}$, $\theta_{\text{A,N}}$, $\tau_{\text{A,L}}$, $\tau_{\text{A,N}}$, $f_{\text{A,L}}$, and $f_{\text{A,N}}$ can be assumed as unchanged within $N_{0}\leq N_{\text{coh}}$ OFDM symbol blocks. Local stationarity of the scattering geometry is widely used in the literature and experimentally confirmed in [46]. Moreover, we consider an underspread A2C channel, characterized by $\max\{\tau_{\text{A,L}},\tau_{\text{A,N}}\}\,f_{\text{max}}\ll 1$, which is true for many wireless channels [47].
[…]
$\sigma_{\text{A,C}}^{2}=\sigma_{\text{A,C,ref}}^{2}\,\mathcal{P}_{\text{A}}\,G_{\text{A}}\,G_{\text{BS,A,C}}\,\left(\frac{d_{\text{A,ref}}}{d_{\text{A,C}}}\right)^{\eta_{\text{A,C}}}$ (8)
[…]
2.4 Receive signal model before CP removal
Hereinafter, without loss of generality, the reported signal models refer to the observation interval $[0,T_{0})$ of the TU channel, which corresponds to the transmission of $N_{0}$ OFDM symbols. After filtering, for $t\in[0,T_{0})$ and $j\in\{1,2,\ldots,J\}$, the baseband received signal at the $j$-th antenna of the BS can be written as
$y_{j}(t)=g_{\text{A,L}}\,e^{\jmath\,2\pi f_{\text{A,L}}t}\,x_{\text{A}}(t-\tau_{\text{A,L}})\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A,L}})}\\ +g_{\text{A,N}}\,e^{\jmath\,2\pi f_{\text{A,N}}t}\,x_{\text{A}}(t-\tau_{\text{A,N}})\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A,N}})}\\ +\sum_{k=1}^{K_{\text{T}}}g_{\text{T},k}\,x_{\text{T}}(t-\tau_{\text{T},k})\,e^{\jmath\pi(j-1)\sin(\theta_{\text{T},k})}+w_{j}(t)$ (9)
where $w_{j}(t)$ is the complex envelope of (filtered) noise statistically independent of $x_{\text{TX}}(t)$, for $\text{TX}\in\{\text{A},\text{T}\}$.
If we denote with $\psi_{\text{ADC}}(t)$ the impulse response of the (anti-aliasing) filter at the input of the analog-to-digital converter (ADC) at the BS, the impulse response of the cascade of the DAC interpolation filter and the ADC antialiasing filter is given by $\psi(t)\triangleq\psi_{\text{DAC}}(t)\ast\psi_{\text{ADC}}(t)$, obeying $\psi(t)\equiv 0$ for $t\not\in[0,L_{\text{filter}}\,T_{\text{c}})$. The pulse $\psi(t)$ and its finite duration $L_{\text{filter}}\,T_{\text{c}}$ are known at the BS. A very common assumption in multicarrier applications is that $L_{\text{filter}}\,T_{\text{c}}+\tau_{\text{max}}<T_{\text{s}}$, with $\tau_{\text{max}}\triangleq\max(\tau_{\text{A,L}},\tau_{\text{A,N}},\tau_{\text{T},1},\ldots,\tau_{\text{T},K_{\text{T}}})$. Moreover, we customarily assume that the BS has been previously locked to the multipath component at (approximately) the minimum delay $\tau_{\text{min}}\triangleq\min(\tau_{\text{A,L}},\tau_{\text{A,N}},\tau_{\text{T},1},\ldots,\tau_{\text{T},K_{\text{T}}})$ and, without loss of generality, we set $\tau_{\text{min}}=0$.
The signal (9) is sampled at $t_{n,p}\triangleq nT_{\text{s}}+p\,T_{\text{c}}$, for $n\in\mathbb{Z}$ and $p\in\{0,1\ldots,P-1\}$. Let $y_{j}^{[p]}[n]\triangleq y_{j}(t_{n,p})$ be the discrete-time counterpart of (9), for $n\in\{0,1,\ldots,N_{0}-1\}$, one gets
$y_{j}^{[p]}[n]=g_{\text{A,L}}\,e^{\jmath\,2\pi\nu_{\text{A,L}}\left(n+\frac{p}{P}\right)}\,x_{\text{A,L}}^{[p]}[n]\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A,L},})}\\ +g_{\text{A,N}}\,e^{\jmath\,2\pi\nu_{\text{A,N}}\left(n+\frac{p}{P}\right)}\,x_{\text{A,N}}^{[p]}[n]\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A,N},})}\\ +\sum_{k=1}^{K_{\text{T}}}g_{\text{T},k}\,x_{\text{T},k}^{[p]}[n]\,e^{\jmath\pi(j-1)\sin(\theta_{\text{T},k})}+w_{j}^{[p]}[n]$ (10)
[…]
$x_{\text{A,C}}^{[p]}[n]$ $=\sum_{\ell=n-1}^{n}\sum_{q=0}^{P-1}u_{\text{A}}^{[q]}[\ell]\,\alpha_{\text{A,C}}\left[(n-\ell)P+(p-q)-d_{\text{A,C}}\right]$ (11)
$x_{\text{T},k}^{[p]}[n]$ $=\sum_{\ell=n-1}^{n}\sum_{q=0}^{P-1}u_{\text{T}}^{[q]}[\ell]\,\alpha_{\text{T},k}\left[(n-\ell)P+(p-q)-d_{\text{T},k}\right]$ (12)
where
$\alpha_{\text{A,C}}[h]$ $\triangleq\psi(h\,T_{\text{c}}-\chi_{\text{A,C}})\quad(h\in\mathbb{Z})$ (13)
$\tau_{\text{A,C}}$ $=d_{\text{A,C}}\,T_{\text{c}}+\chi_{\text{A,C}}$ (14)
$\alpha_{\text{T},k}[h]$ $\triangleq\psi(h\,T_{\text{c}}-\chi_{\text{T},k})$ (15)
$\tau_{\text{T},k}$ $=d_{\text{T},k}\,T_{\text{c}}+\chi_{\text{T},k}$ (16)
with integer delays $d_{\text{A,C}}$ and $d_{\text{T},k}$, and fractional delays $\chi_{\text{A,C}}\in[0,T_{\text{c}})$ and $\chi_{\text{T},k}\in[0,T_{\text{c}})$, and $w_{j}^{[p]}[n]\triangleq w_{j}(t_{n,p})$.
By gathering the obtained samples of the received signal into $\overline{{\mathbf{y}}}_{j}[n]\triangleq(y_{j}^{[0]}[n],y_{j}^{[1]}[n],\ldots,y_{j}^{[P-1]}[n])^{\text{T}}\in\mathbb{C}^{P}$ and accounting for (10), we obtain the following vector model
$\overline{{\mathbf{y}}}_{j}[n]=\overline{{\mathbf{H}}}_{\text{A},j}^{(0)}[n]\,{\mathbf{s}}_{\text{A}}[n]+\overline{{\mathbf{H}}}_{\text{A},j}^{(1)}[n]\,{\mathbf{s}}_{\text{A}}[n-1]\\ +\overline{{\mathbf{H}}}_{\text{T},j}^{(0)}\,{\mathbf{s}}_{\text{T}}[n]+\overline{{\mathbf{H}}}_{\text{T},j}^{(1)}\,{\mathbf{s}}_{\text{T}}[n-1]+\overline{{\mathbf{w}}}_{j}[n]$ (17)
for $n\in\{0,1,\ldots,N_{0}-1\}$ and $j\in\{1,2,\ldots,J\}$, where
$\overline{\mathbf{H}}_{\text{A},j}^{(b)}[n]$ $\triangleq g_{\text{A,L}}\left(\mathbf{D}_{\text{A,L}}\,\mathbf{T}_{\text{A,L}}^{(b)}\,{\mathbf{\Omega}}_{\text{A}}\right)e^{\jmath\,2\pi\nu_{\text{A,L}}n}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A,L}})}$
$+g_{\text{A,N}}\left(\mathbf{D}_{\text{A,N}}\,\mathbf{T}_{\text{A,N}}^{(b)}\,{\mathbf{\Omega}}_{\text{A}}\right)e^{\jmath\,2\pi\nu_{\text{A,N}}n}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A,N}})}$ (18)
$\overline{\mathbf{H}}_{\text{T},j}^{(b)}$ $\triangleq\sum_{k=1}^{K_{\text{T}}}g_{\text{T},k}\left(\mathbf{T}_{\text{T},k}^{(b)}\,{\mathbf{\Omega}}_{\text{T}}\right)e^{\jmath\pi(j-1)\sin(\theta_{\text{T},k})}$ (19)
for $b\in\{0,1\}$, with
$\mathbf{D}_{\text{A,C}}\triangleq{\text{diag}}(1,e^{\jmath\,\frac{2\pi}{P}\nu_{\text{A,C}}},\ldots,e^{\jmath\,\frac{2\pi}{P}\nu_{\text{AC}}(P-1)})$ (20)
[…]
${\mathbf{\Omega}}_{\text{TX}}\triangleq{\mathbf{I}}_{\text{cp}}\,{\mathbf{W}}_{M}\in\mathbb{C}^{P\times M}$ (21)
and $\overline{{\mathbf{w}}}_{j}[n]\triangleq(w^{[0]}_{j}[n],w^{[1]}_{j}[n],\ldots,w^{[P-1]}_{j}[n])^{\text{T}}\in\mathbb{C}^{P}$ is a zero-mean complex circular Gaussian random vector with ${\mathbb{E}}(\overline{{\mathbf{w}}}_{j_{1}}[n_{1}]\,\overline{{\mathbf{w}}}_{j_{2}}^{\text{H}}[n_{2}])=\sigma_{w}^{2}\,\delta_{n_{1}-n_{2}}\,\delta_{j_{1}-j_{2}}\,{\mathbf{I}}_{P}$.
The signal model (17) before CP removal is used in Subsection III-A to blindly estimate the Doppler shifts and the time delays of the AU.
2.5 Receive signal model after CP removal
For $n\in\{0,1,\ldots,N_{0}-1\}$ and $j\in\{1,2,\ldots,J\}$, interblock interference (IBI) in (10) can be perfectly suppressed through CP removal, thus yielding, after DFT,
${\mathbf{y}}_{j}[n]$ $\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\overline{{\mathbf{y}}}_{j}[n]$
$={\mathbf{H}}_{\text{A},j}[n]\,{\mathbf{s}}_{\text{A}}[n]+{\mathbf{H}}_{\text{T},j}\,{\mathbf{s}}_{\text{T}}[n]+{\mathbf{w}}_{j}[n]$ (22)
provided that $L_{\text{cp}}\geq L_{\text{filter}}+\lfloor\tau_{\text{max}}/T_{\text{c}}\rfloor$, where ${\mathbf{R}}_{\text{cp}}\in\mathbb{R}^{M\times P}$ performs CP removal, ${\mathbf{H}}_{\text{A},j}[n]\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\overline{\mathbf{H}}_{\text{A},j}^{(0)}[n]\in\mathbb{C}^{M\times M}$ and ${\mathbf{H}}_{\text{T},j}\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\overline{\mathbf{H}}_{\text{T},j}^{(0)}\in\mathbb{C}^{M\times M}$, and, finally, the noise contribution is ${\mathbf{w}}_{j}[n]\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\overline{{\mathbf{w}}}_{j}[n]\in\mathbb{C}^{M}$. Moreover, accounting for (19), the matrix ${\mathbf{H}}_{\text{T},j}$ can be written as reported at the top of the next page in (23),
${\mathbf{H}}_{\text{T},j}={\mathbf{W}}_{M}^{\text{H}}\sum_{k=1}^{K_{\text{T}}}\underbrace{\left[g_{\text{T},k}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{T},k})}\sum_{\ell=0}^{L_{\text{filter}}}\alpha_{\text{T},k}[\ell]\,{\mathbf{R}}_{\text{cp}}\,{\mathbf{F}}^{\ell+d_{\text{T},k}}\,{\mathbf{I}}_{\text{cp}}\right]}_{{\mathbf{W}}_{M}\,{\mathbf{M}}_{\text{T},j,k}\,{\mathbf{W}}_{M}^{\text{H}}}{\mathbf{W}}_{M}={\mathbf{M}}_{\text{T},j}$ (23)
where the forward matrix ${\mathbf{F}}\in\mathbb{C}^{P\times P}$ has been defined in Subsection I-D and we have observed that the matrix in square brackets is circulant and, thus, it can be equivalently written as ${\mathbf{W}}_{M}\,{\mathbf{M}}_{\text{T},j,k}{\mathbf{W}}_{M}^{\text{H}}$, with
${\mathbf{M}}_{\text{T},j}\triangleq\sum_{k=1}^{K_{\text{T}}}{\mathbf{M}}_{\text{T},j,k}\>.$ (24)
The matrix ${\mathbf{M}}_{\text{T},j,k}\triangleq{\text{diag}}(\boldsymbol{\mu}_{\text{T},j,k})$ is diagonal, with
$\boldsymbol{\mu}_{\text{T},j,k}\triangleq\sqrt{M}\,{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{L}}\,{\mathbf{g}}_{\text{T},j,k}\in\mathbb{C}^{M}$ (25)
where
${\mathbf{L}}$ $\triangleq({\mathbf{I}}_{L_{\text{cp}}},{\mathbf{O}}_{(M-L_{\text{cp}})\times L_{\text{cp}}}^{\text{T}})^{\text{T}}\in\mathbb{R}^{M\times L_{\text{cp}}}$ (26)
${\mathbf{g}}_{\text{T},j,k}$ $\triangleq({\mathbf{0}}_{d_{\text{T},k}}^{\text{T}},g_{\text{T},k}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{T},k})}\,\boldsymbol{\alpha}_{\text{T},k}^{\text{T}},$ (27)
$\hskip 71.13188pt\ldots{\mathbf{0}}_{L_{\text{cp}}-L_{\text{filter}}-d_{\text{T},k}-1}^{\text{T}})^{\text{T}}\in\mathbb{C}^{L_{\text{cp}}}$ (28)
$\boldsymbol{\alpha}_{\text{T},k}$ $\triangleq(\alpha_{\text{T},k}[0],\alpha_{\text{T},k}[1],\ldots,\alpha_{\text{T},k}[L_{\text{filter}}])^{\text{T}}\in\mathbb{R}^{L_{\text{filter}}+1}\>.$ (29)
On the other hand, the transmission of the AU is adversely affected by intercarrier interference (ICI) due to the presence of Doppler shifts: in general, the channel matrix $\mathbf{H}_{\text{A},j}[n]$ cannot be diagonalized through DFT.
By collecting the data received by all the antennas of the BS in the vector ${\mathbf{y}}[n]\triangleq({\mathbf{y}}_{1}^{\text{T}}[n],{\mathbf{y}}^{\text{T}}_{2}[n],\ldots,{\mathbf{y}}^{\text{T}}_{J}[n])^{\text{T}}\in\mathbb{C}^{JM}$, we obtain from (22) and (23) the signal model
${\mathbf{y}}[n]\triangleq{\mathbf{H}}_{\text{A}}[n]\,{\mathbf{s}}_{\text{A}}[n]+{\mathbf{M}}_{\text{T}}\,{\mathbf{s}}_{\text{T}}[n]+{\mathbf{w}}[n]$ (30)
for $n\in\{0,1,\ldots,N_{0}-1\}$, where
${\mathbf{H}}_{\text{A}}[n]$ $\triangleq({\mathbf{H}}_{\text{A},1}^{\text{T}}[n],{\mathbf{H}}_{\text{A},2}^{\text{T}}[n],\ldots,{\mathbf{H}}_{\text{A},J}^{\text{T}}[n])^{\text{T}}\in\mathbb{C}^{(JM)\times M}$ (31)
${\mathbf{M}}_{\text{T}}$ $\triangleq({\mathbf{M}}_{\text{T},1}^{\text{T}},{\mathbf{M}}_{\text{T},2}^{\text{T}},\ldots,{\mathbf{M}}_{\text{T},J}^{\text{T}})^{\text{T}}\in\mathbb{C}^{(JM)\times M}$ (32)
and ${\mathbf{w}}[n]\triangleq({\mathbf{w}}_{1}^{\text{T}}[n],{\mathbf{w}}^{\text{T}}_{2}[n],\ldots,{\mathbf{w}}^{\text{T}}_{J}[n])^{\text{T}}\in\mathbb{C}^{JM}$.
2.6 Pilot allocation scheme
To acquire the channel parameters needed for demodulating the data (i.e., information-bearing) symbols transmitted by the users, the BS might rely on training (or pilot) symbols sent by both the AU and TU, over reserved time-frequency RBs in uplink. Specifically, we assume that user $\text{TX}\in\{\text{A},\text{T}\}$ inserts $1\leq Q_{\text{TX}}\leq M$ pilot symbols at distinct subcarrier locations
$\mathcal{M}_{\text{TX}}\triangleq\{m_{\text{TX},0},m_{\text{TX},1},\ldots,m_{\text{TX},Q_{\text{TX}}-1}\}\subseteq\{0,1,\ldots,M-1\}$ (33)
during the $1\leq N_{\text{TX,train}}\ll N_{0}$ distinct symbol intervals
$\mathcal{N}_{\text{TX}}\triangleq\{n_{\text{TX},1},n_{\text{TX},2},\ldots,n_{\text{TX},N_{\text{TX, train}}}\}\subset\{0,1,\ldots,N_{0}-1\}\>.$ (34)
In NOMA schemes, the training phase of one user suffers from the interference generated by either the pilots or the data transmitted by the other one. Perfect separation of pilot symbols at the BS is possible only if the TU will not transmit neither pilots nor data on the set $\mathcal{M}_{\text{A}}$ of subcarriers over the OFDM signaling intervals $\mathcal{N}_{\text{A}}$ and vice versa. Obviously, besides requiring a stringent protocol-level cooperation between the aerial and terrestrial networks, such a latter strategy comes at a (possibly unaffordable) cost of spectral efficiency due to the time-varying nature of the AU channel and it is not considered herein.
3 Channel estimation for the aerial user
The task of the AU channel estimator is to provide to the WL-TV SIC filter the channel matrices ${\mathbf{H}}_{\text{A},1}[n],{\mathbf{H}}_{\text{A},2}[n],\ldots,{\mathbf{H}}_{\text{A},J}[n]$, for $n\in\{0,1,\ldots,N_{0}-1\}$. In this case, since the time-varying nature of the A2C channel prevents the diagonalization of ${\mathbf{H}}_{\text{A},j}[n]$ through DFT, it is not possible to separate pilots and data symbols in the frequency-domain. For such a reason, we set $\mathcal{M}_{\text{A}}\equiv\{0,1,\ldots,M-1\}$ or, equivalently, $Q_{\text{A}}=M$, i.e., the AU inserts pilots at all the subcarriers within the symbol intervals specified by $\mathcal{N}_{\text{A}}=\{n_{\text{A},1},n_{\text{A},2},\ldots,n_{\text{A},N_{\text{A,train}}}\}$.
To reduce the amount of training, we propose a semi-blind approach to estimate the unknowns characterizing the channel matrices ${\mathbf{H}}_{\text{A},1}[n],{\mathbf{H}}_{\text{A},2}[n],\ldots,{\mathbf{H}}_{\text{A},J}[n]$. Specifically, Doppler and delay parameters are blindly estimated by exploiting the unique ACS properties of the signal transmitted by the AU, without relying on pilots. On the other hand, complex gain and AoA parameters are estimated by using the pilot symbols known by the BS. The estimators derived herein are based on the assumption that the AU channel parameters are deterministic but unknown quantities.
3.1 Blind estimation of Doppler and delay parameters
The starting point of our blind estimation procedure is the signal (17), i.e., the entire OFDM block (including the CP) is processed for acquiring Doppler and delay parameters of the AU transmission. The ACS features of (17) stem from the fact that $\overline{\mathbf{H}}_{\text{A},j}^{(b)}[n]$ in (18) is a discrete-time almost-periodic (AP) matrix [51] with (possibly) incommensurate frequencies $\{\nu_{\text{A,L}},{\nu}_{\text{A,N}}\}\subseteq[-1/2,1/2)$. Consequently, for $r\in\{-1,0,1\}$ and $j\in\{1,2,\ldots,J\}$, the correlation matrix ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}}[n,r]\triangleq{\mathbb{E}}(\overline{{\mathbf{y}}}_{j}[n]\,\overline{{\mathbf{y}}}_{j}^{\text{H}}[n-r])\in\mathbb{C}^{P\times P}$ and its conjugate counterpart ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,r]\triangleq{\mathbb{E}}(\overline{{\mathbf{y}}}_{j}[n]\,\overline{{\mathbf{y}}}_{j}^{\text{T}}[n-r])\in\mathbb{C}^{P\times P}$ are AP matrices, too, and the multivariate process (17) is said to be second-order wide-sense ACS [34].
Capitalizing on the fact that the AU transmits noncircular symbols while the TU adopts a circular modulation format (modulation diversity), the proposed blind estimation approach relies only on the conjugate second-order statistics of the ACS random process (17), which turn out to be free from both TU and noise contributions (a consequence of Doppler diversity). For $j\in\{1,2,\ldots,J\}$, one gets
${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,r]=\sum_{\text{C}_{1}\in\{\text{L, N}\}}\sum_{\text{C}_{2}\in\{\text{L, N}\}}{\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[r]\\ \cdot e^{\jmath 2\pi(\nu_{\text{A},\text{C}_{1}}+\nu_{\text{A},\text{C}_{2}})n}$ (35)
where the matrices in (35) are defined in (36)–(38) at the top of the next page, with ${\mathbf{\Delta}}\in\mathbb{C}^{M\times M}$ being a known diagonal matrix derived in the Appendix for the $\pi/2$-BPSK case and depending on the noncircularity property of the AU symbols.
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[-1]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,e^{\jmath\,2\pi\nu_{\text{A},\text{C}_{2}}}\,(\mathbf{D}_{\text{A},\text{C}_{1}}\,\mathbf{T}_{\text{A},\text{C}_{1}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{D}_{\text{A},\text{C}_{2}}\,\mathbf{T}_{\text{A},\text{C}_{2}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\in\mathbb{C}^{P\times P}$ (36)
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[0]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,\mathbf{D}_{\text{A},\text{C}_{1}}\,[(\mathbf{T}_{\text{A},\text{C}_{1}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A},\text{C}_{2}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$
$\hskip 227.62204pt+(\mathbf{T}_{\text{A},\text{C}_{1}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A},\text{C}_{2}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}]\,\mathbf{D}_{\text{A},\text{C}_{2}}\in\mathbb{C}^{P\times P}$ (37)
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[1]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,e^{-\jmath\,2\pi\nu_{\text{A},\text{C}_{2}}}\,(\mathbf{D}_{\text{A},\text{C}_{1}}\,\mathbf{T}_{\text{A},\text{C}_{1}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{D}_{\text{A},\text{C}_{2}}\,\mathbf{T}_{\text{A},\text{C}_{2}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\in\mathbb{C}^{P\times P}$ (38)
Under mild regularity conditions, the AP matrix (35) can be equivalently written in terms of the generalized Fourier series expansion as follows
${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}^{*}_{j}}[n,r]=\sum_{\alpha}{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]\,e^{\jmath 2\pi\alpha n}$ (39)
where, for $r\in\{-1,0,1\}$,
${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]$ $\triangleq\left\langle{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,r]\,e^{-j2\pi\alpha n}\right\rangle$
$=\lim_{N\to+\infty}\frac{1}{2N+1}\sum_{n=-N}^{N}{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,r]\,e^{-j2\pi\alpha n}$ (40)
is the conjugate cyclic correlation matrix (CCCM) at cycle frequency $\alpha$ and it is given by
${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]=\begin{cases}{\mathbf{\Xi}}_{\text{A},j,\text{L},\text{L}}[r]\>,&\text{$\alpha=2\,\nu_{\text{A,L}}$}\>;\\ {\mathbf{\Xi}}_{\text{A},j,\text{N},\text{N}}[r]\>,&\text{$\alpha=2\,\nu_{\text{A,N}}$}\>;\\ {\mathbf{\Xi}}_{\text{A},j,\text{L},\text{N}}[r]+{\mathbf{\Xi}}_{\text{A},j,\text{N},\text{L}}[r]\>,&\text{$\alpha=\nu_{\text{A,L}}+\nu_{\text{A,N}}$}\>;\\ {\mathbf{O}}_{P\times P}\>,&\text{otherwise}\>.\end{cases}$ (41)
It is worth noticing that there are three distinct cycle frequencies $2\,\nu_{\text{A,L}}$, $2\,\nu_{\text{A,N}}$, and $\nu_{\text{A,L}}+\nu_{\text{A,N}}$, [^6] Since the LoS and NLoS components arrive at the BS from different angular directions, one can safely assume that $\nu_{\text{A,L}}\neq\nu_{\text{A,N}}$. which arise from linear combinations of the Doppler shifts $\nu_{\text{A,L}}$ and $\nu_{\text{A,N}}$. Thus, the cycle frequency set of (35) is
$\mathcal{A}=\left\{2\,\nu_{\text{A,L}},2\,\nu_{\text{A,N}},\nu_{\text{A,L}}+\nu_{\text{A,N}}\right\}\>.$ (42)
Another key observation underlying our blind estimation approach is that, in practice, ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]$ can be directly estimated from the received data using the consistent estimate (see, e.g., [52])
$\widehat{{\mathbf{R}}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]=\frac{1}{N_{0}}\sum_{n=0}^{N_{0}-1}\overline{{\mathbf{y}}}_{j}[n]\,\overline{{\mathbf{y}}}_{j}^{\text{T}}[n-r]\,e^{-j\,2\pi\alpha n}\>.$ (43)
3.1.1 Blind estimation of the Doppler shifts
Estimation of the Doppler shifts $f_{\text{A,L}}$ and $f_{\text{A,N}}$ is equivalent to that of $\nu_{\text{A,L}}$ and $\nu_{\text{A,N}}$, respectively. For $j\in\{1,2,\ldots,J\}$, knowledge of the CCCMs ${\mathbf{R}}_{\overline{{\mathbf{y}}}\overline{{\mathbf{y}}}^{*}}^{\alpha}[-1]$, ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[0]$, and ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[1]$ enables one to blindly retrieve the unknown cycle frequency set (42) through the one-dimensional function:
$J(\alpha)\triangleq\sum_{j=1}^{J}\sum_{r=-1}^{1}\left\|{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]\right\|^{2}\\ =\sum_{j=1}^{J}\sum_{r=-1}^{1}{\text{tr}}\left({\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]\left\{{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{\alpha}[r]\right\}^{\text{H}}\right)\>,\\ \quad\text{with $\alpha\in[-1/2,1/2)$}\>.$ (44)
By using differential calculus arguments, it can be shown that the function $J(\alpha)$ has three local maxima in $[-1/2,1/2)$ at points $2\,\nu_{\text{A,L}}$, $2\,\nu_{\text{A,N}}$, and $\nu_{\text{A,L}}+\nu_{\text{A,N}}$. Hence, the cycle frequencies of the second-order ACS process (17) can be acquired by searching for the maxima of (44) over $[-1/2,1/2)$, provided that $|\nu_{\text{A,L}}|,|\nu_{\text{A,N}}|\leq 1/4$. Specifically, if $\nu_{\text{A,L}}<\nu_{\text{A,N}}$, the maximum points of $J(\alpha)$ in $[-1/2,1/2)$ are $2\,\nu_{\text{A,L}}<\nu_{\text{A,L}}+\nu_{\text{A,N}}<2\,\nu_{\text{A,N}}$ in increasing order. On the contrary, if $\nu_{\text{A,L}}>\nu_{\text{A,N}}$, the increasing order of the maximum points is $2\,\nu_{\text{A,L}}<\nu_{\text{A,L}}+\nu_{\text{A,N}}<2\,\nu_{\text{A,L}}$. Therefore, by picking up the smallest and largest maximum points of $J(\alpha)$ in $[-1/2,1/2)$, one can acquire the Doppler shifts of the AU up to a permutation ambiguity, i.e., it is not possible to infer if the smallest (or, conversely, largest) maximum point of (44) is equal to $2\,\nu_{\text{A,L}}$ or $2\,\nu_{\text{A,N}}$. However, such an ambiguity is irrelevant for the detection process of the symbols transmitted by the AU.
In practice, the estimates of the Doppler shifts can be directly obtained from data by finding the peaks of the function $\widehat{J}(\alpha)$ obtained from (44) by replacing (40) with (43). Consistency and asymptotic normality of $\widehat{J}(\alpha)$ can be proven by paralleling the derivations detailed in [53].
3.1.2 Blind estimation of the time delays
Hereinafter, we assume that the Doppler shifts have been previously acquired as proposed in Subsection III-A-1 and, for the time being, they are assumed perfectly known.
The proposed estimation technique to acquire the delays $\tau_{\text{A,L}}=d_{\text{A,L}}\,T_{\text{c}}+\chi_{\text{A,L}}$ and $\tau_{\text{A,N}}=d_{\text{A,N}}\,T_{\text{c}}+\chi_{\text{A,N}}$ relies again on the AP conjugate correlation matrices ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,-1]$, ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,0]$, and ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}[n,1]$ in (35). In addition, we capitalize on the following parameterization of the Toeplitz matrices $\mathbf{T}_{\text{A,L}}^{(b)}$ and $\mathbf{T}_{\text{A,N}}^{(b)}$ defined in (18) with $b\in\{0,1\}$ in terms of the forward shift ${\mathbf{F}}\in\mathbb{R}^{P\times P}$ and backward shift ${\mathbf{B}}\in\mathbb{R}^{P\times P}$ matrices (see Subsection I-D):
$\mathbf{T}_{\text{A,C}}^{(0)}$ $=\sum_{\ell=0}^{L_{\text{cp}}}\psi(\ell\,T_{\text{c}}-\tau_{\text{A,C}})\,{\mathbf{F}}^{\ell}$ (45)
$\mathbf{T}_{\text{A,C}}^{(1)}$ $=\sum_{\ell=1}^{L_{\text{cp}}}\psi(\ell\,T_{\text{c}}-\tau_{\text{A,C}})\,{\mathbf{B}}^{P-\ell}$ (46)
[…]
${\mathbf{\Phi}}_{\text{A},j,\text{C}}[-1]$ $\triangleq e^{-j2\pi\nu_{\text{A,C}}}\,\mathbf{D}_{\text{A,C}}^{*}\,{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\nu_{\text{A,C}}}[-1]\,\mathbf{D}_{\text{A,C}}^{*}$
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$ (47)
${\mathbf{\Phi}}_{\text{A},j,\text{C}}[0]$ $\triangleq\mathbf{D}_{\text{A,C}}^{*}\,{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\nu_{\text{A,C}}}[0]\,\mathbf{D}_{\text{A,C}}^{*}$
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}[(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$
$\hskip 56.9055pt+(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}]$ (48)
${\mathbf{\Phi}}_{\text{A},j,\text{C}}[1]$ $\triangleq e^{j2\pi\nu_{\text{A,C}}}\,\mathbf{D}_{\text{A,C}}^{*}\,{\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\nu_{\text{A,C}}}[1]\,\mathbf{D}_{\text{A,C}}^{*}$
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\>.$ (49)
The proposed estimation algorithm relies on the fact that
${\mathbf{\Phi}}_{\text{A,C}}$ $\triangleq\sum_{j=1}^{J}\sum_{r=-1}^{1}{\mathbf{\Phi}}_{\text{A},j,\text{C}}[r]$
$=g_{\text{A,C}}^{2}\,{\rm D}_{J}^{*}\left(\sin(\theta_{\text{A,C}})\right)\,\mathbf{C}_{\text{A,C}}\,{\mathbf{\Omega}}_{\text{A}}\,{\mathbf{\Delta}}\,{\mathbf{\Omega}}_{\text{A}}^{\text{T}}\,\mathbf{C}_{\text{A,C}}^{\text{T}}$ (50)
[…]
$\mathbf{c}_{\text{A,C}}\triangleq\Big{(}\psi(-\tau_{\text{A,C}}),\psi(T_{\text{c}}-\tau_{\text{A,C}}),\ldots,\\ \psi(L_{\text{cp}}\,T_{\text{c}}-\tau_{\text{A,C}}),0,\ldots,0\Big{)}^{\text{T}}\>.$ (51)
Such a circulant matrix can be easily diagonalized as
$\mathbf{C}_{\text{A,C}}={\mathbf{W}}_{P}\,{\text{diag}}({\mathbf{v}}_{\text{A,C}})\,{\mathbf{W}}_{P}^{\text{H}}$ (52)
where the $p$th entry of
${\mathbf{v}}_{\text{A,C}}\triangleq\sqrt{P}\,{\mathbf{W}}_{P}^{\text{H}}\,\mathbf{c}_{\text{A,C}}\in\mathbb{C}^{P}$ (53)
is given by $\{{\mathbf{v}}_{\text{A,C}}\}_{p}=\Psi_{\text{A,C}}[p]\,e^{-\jmath\frac{2\pi}{T_{\text{s}}}\tau_{\text{A,C}}\,p}$, with
$\Psi_{\text{A,C}}[p]\approx\begin{cases}\frac{1}{T_{\text{c}}}\,\Psi\left(\frac{p}{T_{\text{s}}}\right)\>,\\ \hskip 28.45274pt\text{for $p\in\left\{0,1,\ldots,\frac{P}{2}-1\right\}$}\\ \frac{1}{T_{\text{c}}}\,\Psi\left(\frac{p-P}{T_{\text{s}}}\right)\,e^{\jmath\frac{2\pi}{T_{\text{c}}}\chi_{\text{A,C}}}\>,\\ \hskip 28.45274pt\text{for $p\in\left\{\frac{P}{2}-1,\frac{P}{2},\ldots,P-1\right\}$}\end{cases}$ (54)
and $\Psi(f)\approx 0$ for $f\not\in(-0.5/T_{\text{c}},0.5/T_{\text{c}})$ being the Fourier transform of $\psi(t)$. Is is worth noticing that $\Psi_{\text{A,C}}[0],\Psi_{\text{A,C}}[1],\ldots,\Psi_{\text{A,C}}[\frac{P}{2}-1]$ are equal to the corresponding coefficients of the $P$-point DFT of $\psi(\ell\,T_{\text{c}})$ and, thus, they are known at the BS.
To acquire the delay $\tau_{\text{A,C}}$, we observe that, by substituting (52) in (50), it results that
${\mathbf{W}}_{P}^{\text{H}}\,{\mathbf{\Phi}}_{\text{A,C}}\,{\mathbf{W}}_{P}^{*}=g_{\text{A,C}}^{2}\,{\rm D}_{J}^{*}\left(\sin(\theta_{\text{A,C}})\right)\\ \cdot{\mathbf{E}}_{\text{A,C}}\,{\mathbf{\Psi}}_{\text{A,C}}\,{\mathbf{\Upsilon}}\,{\mathbf{\Psi}}_{\text{A,C}}\,{\mathbf{E}}_{\text{A,C}}$ (55)
[…]
${\mathbf{E}}_{\text{A,C}}$ $\triangleq{\text{diag}}\left(1,e^{-\jmath\frac{2\pi}{T_{\text{s}}}\tau_{\text{A,C}}},\ldots,e^{-\jmath\frac{2\pi}{T_{\text{s}}}\tau_{\text{A,C}}(P-1)}\right)$ (56)
${\mathbf{\Psi}}_{\text{A,C}}$ $\triangleq{\text{diag}}\left(\Psi_{\text{A,C}}[0],\Psi_{\text{A,C}}[1],\ldots,\Psi_{\text{A,C}}[P-1]\right)$ (57)
${\mathbf{\Upsilon}}$ $\triangleq{\mathbf{W}}_{P}^{\text{H}}\,{\mathbf{\Omega}}_{\text{A}}\,{\mathbf{\Delta}}\,{\mathbf{\Omega}}_{\text{A}}^{\text{T}}\,{\mathbf{W}}_{P}^{*}\in\mathbb{C}^{P\times P}\>.$ (58)
For $p\in\{0,1,\ldots,P-1\}$, it is readily seen that the $p$th diagonal entry of ${\mathbf{W}}_{P}^{\text{H}}\,{\mathbf{\Phi}}_{\text{A,C}}\,{\mathbf{W}}_{P}^{*}$ is given by
$\{{\mathbf{W}}_{P}^{\text{H}}\,{\mathbf{\Phi}}_{\text{A,C}}\,{\mathbf{W}}_{P}^{*}\}_{p,p}\\ =g_{\text{A,C}}^{2}\,{\rm D}_{J}^{*}\left(\sin(\theta_{\text{A,C}})\right)\Psi_{\text{A,C}}^{2}[p]\,\{{\mathbf{\Upsilon}}\}_{p,p}\,e^{-j\frac{4\pi}{T_{\text{s}}}\tau_{\text{A,C}}\,p}\>.$ (59)
At this point, let us introduce the one-dimensional cost function defined in (60) at the top of the next page.
$\mathcal{I}_{\text{A,C}}(\beta)$ $\triangleq\left|\sum_{p=0}^{P/2-1}\{{\mathbf{W}}_{P}^{\text{H}}\,{\mathbf{\Phi}}_{\text{A,C}}\,{\mathbf{W}}_{P}^{*}\}_{p,p}\,(\Psi_{\text{A,C}}^{2}[p])^{*}\,\{{\mathbf{\Upsilon}}\}_{p,p}^{*}\,e^{j\frac{4\pi}{T_{\text{s}}}\beta\,p}\right|$
$=|g_{\text{A,C}}|^{2}\,\left|{\rm D}_{J}^{*}\left(\sin(\theta_{\text{A,C}})\right)\right|\,\left|\sum_{p=0}^{P/2-1}\left|\Psi_{\text{A,C}}[p]\right|^{4}\,\left|\{{\mathbf{\Upsilon}}\}_{p,p}\right|^{2}e^{-j\frac{4\pi}{T}(\tau_{\text{A,C}}-\beta)\,p}\right|\>,\quad\text{with $\beta\in[0,\Delta_{\text{max}}]$}$ (60)
[…]
In practice, an estimate $\widehat{\tau}_{\text{A,C}}$ of the delay $\tau_{\text{A,C}}$ is derived by searching for the peak of the finite-sample version ^$\widehat{\mathcal{I}}_{\text{A,C}}(\beta)$ of the cost function $\mathcal{I}_{\text{A,C}}(\beta)$ defined in (60), where ^$\widehat{\mathcal{I}}_{\text{A,C}}(\beta)$ is obtained by replacing ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\nu_{\text{A,C}}}[r]$ in (47)–(49), for $r\in\{-1,0,1\}$, with their corresponding estimates $\widehat{{\mathbf{R}}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\widehat{\nu}_{\text{A,C}}}[r]$ [see (43)], evaluated at the estimated cycle frequency $2\,\widehat{\nu}_{\text{A,C}}$ acquired through the algorithm in Subsection II-A-1. The performance of the proposed estimator for time delays is assessed by simulations in Section 5.
3.2 Pilot-aided estimation of complex gain and AoA parameters
After acquiring the Doppler and delay parameters blindly, we propose to estimate the complex gain and AoA unknowns of the AU, by capitalizing on its pilots transmitted in the symbol training intervals $\mathcal{N}_{\text{A}}=\{n_{\text{A},1},n_{\text{A},2},\ldots,n_{\text{A},N_{\text{A,train}}}\}$. For the present, Doppler and delay parameters are assumed to be perfectly known. The starting point of the pilot-aided estimation step is the signal model (22), which collects the signal received by the $j$-th antenna after CP removal and DFT.
For $n\in\mathcal{N}_{\text{A}}$ and $j\in\{1,2,\ldots,J\}$, it is convenient to rewrite (22) as follows
${\mathbf{y}}_{j}[n]=g_{\text{A,L}}\,\eta_{\text{A,L}}^{j-1}\,\mathbf{p}_{\text{A,L}}[n]\\ +g_{\text{A,N}}\,\eta_{\text{A,N}}^{j-1}\,\mathbf{p}_{\text{A,N}}[n]+{\mathbf{d}}_{\text{A},j}[n]$ (61)
[…]
Given the AU symbol vector ${\mathbf{s}}_{\text{A}}[n]$, the random vector ${\mathbf{y}}_{j}[n]$ is circular, for $j\in\{1,2,\ldots,J\}$. By gathering all the training blocks received by the $j$-th antenna in the matrix ${\mathbf{Y}}_{\text{A},j}\triangleq({\mathbf{y}}_{j}[n_{\text{A},1}],{\mathbf{y}}_{j}[n_{\text{A},2}],\ldots,{\mathbf{y}}_{j}[n_{\text{A},N_{\text{A,train}}}])\in\mathbb{C}^{M\times N_{\text{A,train}}}$, one gets the compact matrix model
${\mathbf{Y}}_{\text{A},j}=g_{\text{A,L}}\,\eta_{\text{A,L}}^{j-1}\,\mathbf{P}_{\text{A,L}}+g_{\text{A,N}}\,\eta_{\text{A,N}}^{j-1}\,\mathbf{P}_{\text{A,N}}+{\mathbf{D}}_{\text{A},j}$ (62)
with
$\mathbf{P}_{\text{A,C}}$ $\triangleq(\mathbf{p}_{\text{A,C}}[n_{\text{A},1}],\mathbf{p}_{\text{A,C}}[n_{\text{A},2}],$
$\hskip 56.9055pt\ldots,\mathbf{p}_{\text{A,C}}[n_{\text{A},N_{\text{A,train}}}])\in\mathbb{C}^{M\times N_{\text{A,train}}}$ (63)
${\mathbf{D}}_{\text{A},j}$ $\triangleq({\mathbf{d}}_{\text{A},j}[n_{\text{A},1}],{\mathbf{d}}_{\text{A},j}[n_{\text{A},2}],$
$\hskip 56.9055pt\ldots,{\mathbf{d}}_{\text{A},j}[n_{\text{A},N_{\text{A,train}}}])\in\mathbb{C}^{M\times N_{\text{A,train}}}$ (64)
[…]
Since the correlation properties of ${\mathbf{D}}_{\text{A},j}$ are unknown due to the multiple access interference arising from the transmission of the TU, estimation strategies involving the second-order statistics of the disturbance, such as best linear unbiased (BLU) or linear MMSE estimators, cannot be implemented in this case. Therefore, we resort to the least-squares (LS) criterion [54] to acquire channel gains and AoAs of the AU, whose salient feature is that no probability assumptions are made about the received data.
The four parameters $\eta_{\text{A,L}}$, $g_{\text{A,L}}$, $\eta_{\text{A,N}}$, and $g_{\text{A,N}}$ have to be estimated. The LS estimator is defined as
[…]
with
${\mathcal{L}}_{A}(\rho_{\text{L}},\rho_{\text{N}},\varrho_{\text{L}},\varrho_{\text{N}})\triangleq\sum_{j=1}^{J}\left\|{\mathbf{Y}}_{\text{A},j}-\sum_{\text{C}\in\{\text{L, N}\}}\rho_{\text{C}}\,\,\varrho_{\text{C}}^{j-1}\,\mathbf{P}_{\text{A,C}}\right\|^{2}\>.$ (66)
Significant computational savings follow from the observation that, for any given $\eta_{\text{A,L}}$ and $\eta_{\text{A,N}}$, the problem of finding the LS estimates of $g_{\text{A,L}}$ and $g_{\text{A,N}}$, corresponding to that $\eta_{\text{A,L}}$ and $\eta_{\text{A,N}}$, are obtained by solving the $2\times 2$ linear system:
$\frac{\partial}{\partial\rho_{\text{L}}^{*}}\,{\mathcal{L}}_{A}(\rho_{\text{L}},\rho_{\text{N}},\varrho_{\text{L}},\varrho_{\text{N}})$ $=-\Lambda_{\text{L},0}+\Lambda_{\text{L},1}\,\rho_{\text{L}}+\Lambda_{\text{L},2}\,\rho_{\text{N}}=0$ (67)
$\frac{\partial}{\partial\rho_{\text{N}}^{*}}\,{\mathcal{L}}_{A}(\rho_{\text{L}},\rho_{\text{N}},\varrho_{\text{L}},\varrho_{\text{N}})$ $=-\Lambda_{\text{N},0}+\Lambda_{\text{N},1}\,\rho_{\text{N}}+\Lambda_{\text{N},2}\,\rho_{\text{L}}=0$ (68)
with
$\Lambda_{\text{C},0}$ $\triangleq\sum_{j=1}^{J}{\text{tr}}\left({\mathbf{Y}}_{\text{A},j}\,\mathbf{P}_{\text{A,C}}^{\text{H}}\,\,\eta_{\text{A,C}}^{-(j-1)}\right)$ (69)
$\Lambda_{\text{C},1}$ $\triangleq J\,{\text{tr}}(\mathbf{P}_{\text{A,C}}\,\mathbf{P}_{\text{A,C}}^{\text{H}})=J\,\|\mathbf{P}_{\text{A,C}}\|^{2}$ (70)
$\Lambda_{\text{C},2}$ $\triangleq\sum_{j=1}^{J}{\text{tr}}\left(\mathbf{P}_{\text{A},\overline{\text{C}}}\,\mathbf{P}_{\text{A,C}}^{\text{H}}\,(\eta_{\text{A},\overline{\text{C}}}\,\eta_{\text{A,C}}^{*})^{j-1}\right)$ (71)
[…]
$\widehat{g}_{\text{A,L}}$ $=\frac{\Lambda_{\text{L},0}\,\Lambda_{\text{N},1}-\Lambda_{\text{L},2}\,\Lambda_{\text{N},0}}{\Lambda_{\text{L},1}\,\Lambda_{\text{N},1}-\Lambda_{\text{L},2}\,\Lambda_{\text{N},2}}$ (72)
$\widehat{g}_{\text{A,N}}$ $=\frac{\Lambda_{\text{L},1}\,\Lambda_{\text{N},0}-\Lambda_{\text{L},0}\,\Lambda_{\text{N},2}}{\Lambda_{\text{L},1}\,\Lambda_{\text{N},1}-\Lambda_{\text{L},2}\,\Lambda_{\text{N},2}}\>.$ (73)
By substituting (72)-(73) into (65), the LS estimates of $\eta_{\text{A,L}}$ and $\eta_{\text{A,N}}$ can be found by
$(\widehat{\eta}_{\text{A,L}},\widehat{\eta}_{\text{A,N}})=\arg\min_{|\varrho_{\text{L}}|=|\varrho_{\text{N}}|=1}{\mathcal{L}}_{A}(\widehat{g}_{\text{A,L}},\widehat{g}_{\text{A,N}},\varrho_{\text{L}},\varrho_{\text{C}})\>.$ (74)
In general, the cost function in (74) is multimodal. However, there exist computationally-efficient global optimization procedures, which try to provide an efficiently global optimization by partition of definition space at multi-scale levels, see, e.g., [55, 56]. Such methods have a stable convergence speed and the optimization solution is independent of the selection of initial solutions [55].
In practice, the previously obtained estimates of the Doppler and delay parameters (see Subsection III-A-1 and III-A-2) have to be used for LS acquisition of the complex gain and AoA unknowns of the AU. Specifically, the LS problem to be solved comes from replacing $\mathbf{P}_{\text{A,C}}$ in (66) with its corresponding estimate $\widehat{\mathbf{P}}_{\text{A,C}}\triangleq(\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},1}],\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},2}],\ldots,\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},N_{\text{A,train}}}])$, where $\widehat{\mathbf{p}}_{\text{A,C}}[n]\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\widehat{\mathbf{D}}_{\text{A,C}}\,\widehat{\mathbf{T}}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}}\,{\mathbf{s}}_{\text{A}}[n]\,e^{\jmath\,2\pi\widehat{\nu}_{\text{A,C}}n}$, for any $n\in\mathcal{N}_{\text{A}}$, and $\widehat{\mathbf{T}}_{\text{A,C}}^{(0)}$ is built from (45) by using $\widehat{\tau}_{\text{A,C}}$ in lieu of $\tau_{\text{A,C}}$. The performance of this LS estimator is assessed by simulations in Section 5.
4 Channel estimation for the terrestrial user
The estimation of the AU channel follows an array signal processing approach, where the parameters characterizing the AU channel are estimated. We resort to a completely different approach for the estimation of the TU channel. More precisely, the TU channel estimator passes to the SIC detector the diagonal time-invariant matrix ${\mathbf{H}}_{\text{T},j}\equiv{\mathbf{M}}_{\text{T},j}$, for each antenna index $j\in\{1,2,\ldots,J\}$. Since the TU channel is time invariant, we pursue a pilot-aided approach for estimating ${\mathbf{M}}_{\text{T},j}$, which relies on the $Q_{\text{T}}$ pilot symbols transmitted by the TU at the subcarriers $\mathcal{M}_{\text{T}}=\{m_{\text{T},0},m_{\text{T},1},\ldots,m_{\text{T},Q_{\text{T}}-1}\}$ during the $N_{\text{T,train}}$ symbol intervals $\mathcal{N}_{\text{T}}=\{n_{\text{T},1},n_{\text{T},2},\ldots,n_{\text{T},N_{\text{T,train}}}\}$. This task is developed by assuming perfect knowledge of the AU channel matrices ${\mathbf{H}}_{\text{A},1}[n],{\mathbf{H}}_{\text{A},2}[n],\ldots,{\mathbf{H}}_{\text{A},J}[n]$, which can be acquired as proposed in Section 3. To this aim, according to (25), instead of separately estimating the parameters (i.e., time delays, AoAs, and channel gains) characterizing the TU channel, we choose to directly estimate the vector
${\mathbf{g}}_{\text{T},j}\triangleq\sum_{k=1}^{K_{\text{T}}}{\mathbf{g}}_{\text{T},j,k}\in\mathbb{C}^{L_{\text{cp}}}$ (75)
which gathers all the aggregated state information of the TU channel corresponding to the $j$th receiving antenna.
Let us consider the frequency-domain received data given by (22). Denoting by
${\mathbf{y}}_{\text{T},j}[n]$ $\triangleq(y_{j}^{[L_{\text{cp}}+m_{\text{T},0}]}[n],y_{j}^{[L_{\text{cp}}+m_{\text{T},1}]}[n],$
$\hskip 56.9055pt\ldots,y_{j}^{[L_{\text{cp}}+m_{\text{T},Q_{\text{T}}-1}]}[n])^{\text{T}}\in\mathbb{C}^{Q_{\text{T}}}$ (76)
${\mathbf{w}}_{\text{T},j}[n]$ $\triangleq(w_{j}^{[L_{\text{cp}}+m_{\text{T},0}]}[n],w_{j}^{[L_{\text{cp}}+m_{\text{T},1}]}[n],$
$\hskip 56.9055pt\ldots,w_{j}^{[L_{\text{cp}}+m_{\text{T},Q_{\text{T}}-1}]}[n])^{\text{T}}\in\mathbb{C}^{Q_{\text{T}}}$ (77)
the vectors containing the entries of ${\mathbf{y}}_{j}[n]$ and ${\mathbf{w}}_{j}[n]$ at the TU pilot locations, respectively, we get
${\mathbf{y}}_{\text{T},j}[n]=\mathbf{P}_{\text{T}}[n]\,{\mathbf{g}}_{\text{T},j}+{\mathbf{d}}_{\text{T},j}[n]$ (78)
for $n\in\mathcal{N}_{\text{T}}$ and $j\in\{1,2,\ldots,J\}$, where
$\mathbf{P}_{\text{T}}[n]\triangleq\sqrt{M}\,{\mathbf{S}}_{\text{T}}([n])\,{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{L}}\in\mathbb{C}^{Q_{\text{T}}\times L_{\text{cp}}}$ (79)
is a known matrix, with the pilot matrix ${\mathbf{S}}_{\text{T}}([n])\in\mathbb{C}^{Q_{\text{T}}\times M}$ obtained by picking up the rows of ${\text{diag}}({\mathbf{s}}_{\text{T}}[n])$ at positions $\{m_{\text{T},0}+1,m_{\text{T},1}+1,\ldots,m_{\text{T},Q_{\text{T}}-1}+1\}$ for $n\in\mathcal{N}_{\text{T}}$, whereas ${\mathbf{d}}_{\text{T},j}[n]\triangleq{\mathbf{M}}_{\text{A},j}[n]\,{\mathbf{s}}_{\text{A}}[n]+{\mathbf{w}}_{j}[n]$ represents disturbance from the TU viewpoint, with ${\mathbf{M}}_{\text{A},j}[n]\in\mathbb{C}^{Q_{\text{T}}\times M}$ obtained by taking the rows of the known matrix ${\mathbf{H}}_{\text{A},j}[n]$ corresponding to the indexes $\{m_{\text{T},0}+1,m_{\text{T},1}+1,\ldots,m_{\text{T},Q_{\text{T}}-1}+1\}$. The training blocks received by the $j$-th antenna for $n\in\mathcal{N}_{\text{T}}$ are gathered into the $(Q_{\text{T}}N_{\text{T,train}})$-dimensional complex vector
${\mathbf{y}}_{\text{T},j}\triangleq({\mathbf{y}}_{\text{T},j}^{\text{T}}[n_{\text{T},1}],{\mathbf{y}}_{\text{T},j}^{\text{T}}[n_{\text{T},2}],\ldots,{\mathbf{y}}_{\text{T},j}^{\text{T}}[n_{\text{T},N_{\text{T,train}}}])^{\text{T}}$ (80)
thus yielding, for $j\in\{1,2,\ldots,J\}$,
${\mathbf{y}}_{\text{T},j}=\mathbf{P}_{\text{T}}\,{\mathbf{g}}_{\text{T},j}+{\mathbf{d}}_{\text{T},j}$ (81)
where
$\mathbf{P}_{\text{T}}$ $\triangleq(\mathbf{P}_{\text{T}}^{\text{T}}[n_{\text{T},1}],\mathbf{P}_{\text{T}}^{\text{T}}[n_{\text{T},2}],$
$\hskip 56.9055pt\ldots\mathbf{P}_{\text{T}}^{\text{T}}[n_{\text{T},N_{\text{T,train}}}])^{\text{T}}\in\mathbb{C}^{(Q_{\text{T}}N_{\text{T,train}})\times L_{\text{cp}}}$ (82)
${\mathbf{d}}_{\text{T},j}$ $\triangleq({\mathbf{d}}_{\text{T},j}^{\text{T}}[n_{\text{T},1}],{\mathbf{d}}_{\text{T},j}^{\text{T}}[n_{\text{T},2}],$
$\hskip 56.9055pt\ldots,{\mathbf{d}}_{\text{T},j}^{\text{T}}[n_{\text{T},N_{\text{T,train}}}])^{\text{T}}\in\mathbb{C}^{Q_{\text{T}}N_{\text{T,train}}}\>.$ (83)
Finally, we collect the data received by all the antennas of the BS into ${\mathbf{y}}_{\text{T}}\triangleq({\mathbf{y}}_{\text{T},1}^{\text{T}},{\mathbf{y}}_{\text{T},2}^{\text{T}},\ldots,{\mathbf{y}}_{\text{T},J}^{\text{T}})^{\text{T}}\in\mathbb{C}^{JQ_{\text{T}}N_{\text{T,train}}}$, thus yielding the vector signal model
${\mathbf{y}}_{\text{T}}=\left({\mathbf{I}}_{J}\otimes\mathbf{P}_{\text{T}}\right){\mathbf{g}}_{\text{T}}+{\mathbf{d}}_{\text{T}}$ (84)
where
${\mathbf{g}}_{\text{T}}$ $\triangleq({\mathbf{g}}_{\text{T},1}^{\text{T}},{\mathbf{g}}_{\text{T},2}^{\text{T}},\ldots,{\mathbf{g}}_{\text{T},J}^{\text{T}})^{\text{T}}\in\mathbb{C}^{JL_{\text{cp}}}$ (85)
${\mathbf{d}}_{\text{T}}$ $\triangleq({\mathbf{d}}_{\text{T},1}^{\text{T}},{\mathbf{d}}_{\text{T},2}^{\text{T}},\ldots,{\mathbf{d}}_{\text{T},J}^{\text{T}})^{\text{T}}\in\mathbb{C}^{JQ_{\text{T}}N_{\text{T,train}}}\;.$ (86)
We assume that $\mathbf{P}_{\text{T}}$ is full-column rank, i.e., ${\text{rank}}(\mathbf{P}_{\text{T}})=L_{\text{cp}}$ and $Q_{\text{T}}N_{\text{T,train}}\geq L_{\text{cp}}$, which ensures that ${\mathbf{g}}_{\text{T}}$ is uniquely identifiable from (84) in the absence of the disturbance ${\mathbf{d}}_{\text{T}}$.
There are different estimation strategies that can be pursued to estimate the TU channel from (84). We focus on estimators belonging to the family of the classical estimation approaches [54], for which ${\mathbf{g}}_{\text{T}}$ is viewed as deterministic but unknown vector. The simplest estimator is represented by the LS one (see, e.g., [54]):
$\widehat{{\mathbf{g}}}_{\text{T},\text{LS}}$ $\triangleq\arg\min_{{\mathbf{g}}_{\text{T}}\in\mathbb{C}^{JL_{\text{cp}}}}\|{\mathbf{y}}_{\text{T}}-\mathbf{P}_{\text{T}}\,{\mathbf{g}}_{\text{T}}\|^{2}$
$=\left[{\mathbf{I}}_{J}\otimes\left(\mathbf{P}_{\text{T}}^{\text{H}}\,\mathbf{P}_{\text{T}}\right)^{-1}\mathbf{P}_{\text{T}}^{\text{H}}\right]{\mathbf{y}}_{\text{T}}$ (87)
which does not exploit the CSI knowledge of the AU and, thus, performs poorly in a NOMA setting (see Section 5). On the other hand, the BS can jointly exploits the knowledge of the AU channel matrices ${\mathbf{H}}_{\text{A},1}[n],{\mathbf{H}}_{\text{A},2}[n],\ldots,{\mathbf{H}}_{\text{A},J}[n]$ and the noncircularity of the AU symbols, by resorting to the best WL unbiased (BWLU) estimator, which is restricted to be WL in the data, it is unbiased, and it has minimum variance. In matrix form, a WL estimator is defined by
$\widehat{{\mathbf{g}}}_{\text{T}}$ $={\mathbf{A}}_{1}\,{\mathbf{y}}_{\text{T}}+{\mathbf{A}}_{2}\,{\mathbf{y}}_{\text{T}}^{*}$
$=\underbrace{({\mathbf{A}}_{1},{\mathbf{A}}_{2})}_{{\mathbf{A}}\in\mathbb{C}^{(JL_{\text{cp}})\times(2JQ_{\text{T}}N_{\text{T,train}})}}\,\underbrace{\begin{pmatrix}{\mathbf{y}}_{\text{T}}\\ {\mathbf{y}}_{\text{T}}^{*}\end{pmatrix}}_{\widetilde{{\mathbf{y}}}_{\text{T}}\in\mathbb{C}^{2JQ_{\text{T}}N_{\text{T,train}}}}={\mathbf{A}}\,\widetilde{{\mathbf{y}}}_{\text{T}}$ (90)
where ${\mathbf{A}}_{1},{\mathbf{A}}_{2}\in\mathbb{C}^{(JL_{\text{cp}})\times(JQ_{\text{T}}N_{\text{T,train}})}$ and
$\widetilde{{\mathbf{y}}}_{\text{T}}=\widetilde{\mathbf{P}}_{\text{T}}\,{\mathbf{g}}_{\text{T}}+{\mathbf{J}}_{JQ_{\text{T}}N_{\text{T,train}}}\,\widetilde{\mathbf{P}}_{\text{T}}^{*}\,{\mathbf{g}}_{\text{T}}^{*}+\widetilde{{\mathbf{d}}}_{\text{T}}$ (91)
with
$\widetilde{\mathbf{P}}_{\text{T}}\triangleq({\mathbf{I}}_{J}\otimes\mathbf{P}_{\text{T}}^{\text{T}},{\mathbf{O}}_{(JQ_{\text{T}}N_{\text{T,train}})\times(JL_{\text{cp}})}^{\text{T}})^{\text{T}}\in\mathbb{C}^{(2JQ_{\text{T}}N_{\text{T,train}})\times(JL_{\text{cp}})}$ (92)
${\mathbf{J}}_{JQ_{\text{T}}N_{\text{T,train}}}\in\mathbb{R}^{(2JQ_{\text{T}}N_{\text{T,train}})\times(2JQ_{\text{T}}N_{\text{T,train}})}$ defined in (1), and, finally, $\widetilde{{\mathbf{d}}}_{\text{T}}\triangleq({\mathbf{d}}_{\text{T}}^{\text{T}},{\mathbf{d}}_{\text{T}}^{\text{H}})^{\text{T}}\in\mathbb{C}^{2JQ_{\text{T}}N_{\text{T,train}}}$. It is readily verified that $\widetilde{{\mathbf{d}}}_{\text{T}}$ has zero mean, i.e., ${\mathbb{E}}(\widetilde{{\mathbf{d}}}_{\text{T}})={\mathbf{0}}_{2JQ_{\text{T}}N_{\text{T,train}}}$, and
${\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}\triangleq{\mathbb{E}}(\widetilde{{\mathbf{d}}}_{\text{T}}\,\widetilde{{\mathbf{d}}}_{\text{T}}^{\text{H}})=\widetilde{{\mathbf{M}}}_{\text{A}}\,\widetilde{{\mathbf{M}}}_{\text{A}}^{\text{H}}+\sigma_{w}^{2}\,{\mathbf{I}}_{2JQ_{\text{T}}N_{\text{T,train}}}$ (93)
where
$\widetilde{{\mathbf{M}}}_{\text{A}}\triangleq\begin{pmatrix}{\mathbf{M}}_{\text{A}}\\ {\mathbf{M}}_{\text{A}}^{*}\,[{\mathbf{I}}_{N_{\text{T,train}}}\otimes\mathbf{\Delta}]\end{pmatrix}\in\mathbb{C}^{(2JQ_{\text{T}}N_{\text{T,train}})\times(MN_{\text{T,train}})}$ (94)
with
${\mathbf{M}}_{\text{A}}$ $\triangleq({\mathbf{M}}_{\text{A},1}^{\text{T}},{\mathbf{M}}_{\text{A},2}^{\text{T}},\ldots,{\mathbf{M}}_{\text{A},J}^{\text{T}})^{\text{T}}\in\mathbb{C}^{(JQ_{\text{T}}N_{\text{T,train}})\times(MN_{\text{T,train}})}$
${\mathbf{M}}_{\text{A},j}$ $\triangleq{\text{diag}}({\mathbf{M}}_{\text{A},j}[n_{\text{T},1}],{\mathbf{M}}_{\text{A},j}[n_{\text{T},2}],\ldots,{\mathbf{M}}_{\text{A},j}[n_{\text{T},N_{\text{T,train}}}])\>.$ (95)
The key point of the proposed BWLU estimation approach is the fact that the correlation matrix ${\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}$ is known at the BS after acquiring the AU channel state (see Section 3).
According to (91), in order for $\widehat{{\mathbf{g}}}_{\text{T}}$ to be unbiased, we require ${\mathbb{E}}(\widehat{{\mathbf{g}}}_{\text{T}})={\mathbf{A}}\,{\mathbb{E}}(\widetilde{{\mathbf{y}}}_{\text{T}})={\mathbf{g}}_{\text{T}}$, which is tantamount to imposing the unbiased constraint
${\mathbf{A}}\underbrace{\begin{pmatrix}{\mathbf{I}}_{J}\otimes\mathbf{P}_{\text{T}}&\mathbf{O}_{(JQ_{\text{T}}N_{\text{T,train}})\times(JL_{\text{cp}})}\\ \mathbf{O}_{(JQ_{\text{T}}N_{\text{T,train}})\times(JL_{\text{cp}})}&{\mathbf{I}}_{J}\otimes\mathbf{P}_{\text{T}}^{*}\end{pmatrix}}_{{\mathbf{\Pi}}_{\text{T}}\in\mathbb{C}^{(2JQ_{\text{T}}N_{\text{T,train}})\times(2JL_{\text{cp}})}}\\ =\underbrace{\left({\mathbf{I}}_{JL_{\text{cp}}},\mathbf{O}_{(JL_{\text{cp}})\times(JL_{\text{cp}})}\right)}_{{\mathbf{\Theta}}\in\mathbb{R}^{(JL_{\text{cp}})\times(2JL_{\text{cp}})}}$ (96)
The BWLU estimator is found by minimizing the variance at the estimator output
$\text{Var}(\widehat{{\mathbf{g}}}_{\text{T}})$ $\triangleq{\mathbb{E}}\left(\left\|\widehat{{\mathbf{g}}}_{\text{T}}-{\mathbf{g}}_{\text{T}}\right\|^{2}\right)$
$={\text{tr}}\left[{\mathbf{A}}\,{\mathbb{E}}(\widetilde{{\mathbf{y}}}_{\text{T}}\,\widetilde{{\mathbf{y}}}_{\text{T}}^{H})\,{\mathbf{A}}^{\text{H}}\right]-\|{\mathbf{g}}_{\text{T}}\|^{2}$
$={\text{tr}}({\mathbf{A}}\,{\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}\,{\mathbf{A}}^{\text{H}})$ (97)
where we have also taken into account the constraint (96). The classical approach for minimizing (97) subject to the unbiased constraint (96) is the method of Lagrange multipliers, which yields the solution (see, e.g., [54])
$\widehat{{\mathbf{g}}}_{\text{T},\text{BWLU}}={\mathbf{\Theta}}\left({\mathbf{\Pi}}_{\text{T}}^{\text{H}}\,{\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}^{-1}{\mathbf{\Pi}}_{\text{T}}\right)^{-1}{\mathbf{\Pi}}_{\text{T}}^{\text{H}}\,{\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}^{-1}\widetilde{{\mathbf{y}}}_{\text{T}}$ (98)
whose corresponding minimum variance is given by
$\text{Var}(\widehat{{\mathbf{g}}}_{\text{T},\text{BWLU}})={\text{tr}}\left[{\mathbf{\Theta}}\left({\mathbf{\Pi}}_{\text{T}}^{\text{H}}\,{\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}^{-1}{\mathbf{\Pi}}_{\text{T}}\right)^{-1}{\mathbf{\Theta}}^{\text{H}}\right]\>.$ (99)
It can be verified by direct inspection that
${\mathbf{A}}_{\text{BWLU}}$ $={\mathbf{\Theta}}\left({\mathbf{\Pi}}_{\text{T}}^{\text{H}}\,{\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}^{-1}{\mathbf{\Pi}}_{\text{T}}\right)^{-1}{\mathbf{\Pi}}_{\text{T}}^{\text{H}}\,{\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}^{-1}$
$=({\mathbf{A}}_{\text{BWLU},1},{\mathbf{A}}_{\text{BWLU},2})$ (100)
with ${\mathbf{A}}_{\text{BWLU},2}\neq\mathbf{O}_{(JL_{\text{cp}})\times(JQ_{\text{T}}N_{\text{T,train}})}$: in this case, the BWLU estimator is expected to outperform its BLU counterpart in the minimum-variance sense.
In practice, the BWLU estimator is built by using estimates $\widehat{{\mathbf{H}}}_{\text{A},1}[n],\widehat{{\mathbf{H}}}_{\text{A},2}[n],\ldots,\widehat{{\mathbf{H}}}_{\text{A},J}[n]$ of the AU channel matrices ${\mathbf{H}}_{\text{A},1}[n],{\mathbf{H}}_{\text{A},2}[n],\ldots,{\mathbf{H}}_{\text{A},J}[n]$, which are obtained through the three-step approach proposed in Section 3. Such an estimator is implemented by replacing ${\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}$ in (98) with its corresponding estimate $\widehat{{\mathbf{R}}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}$ coming from $\widehat{{\mathbf{H}}}_{\text{A},1}[n],\widehat{{\mathbf{H}}}_{\text{A},2}[n],\ldots,\widehat{{\mathbf{H}}}_{\text{A},J}[n]$, for $n\in\mathcal{N}_{\text{T}}$.
5 Numerical results
In this section, we provide numerical results aimed at evaluating the performance of the proposed channel estimators for the SG-NOMA scheme. To this aim, we consider the following simulation setting. The AU and TU employ OFDM modulation with $M=16$ subcarriers, CP of length $L_{\text{cp}}=4$, sampling rate $6251/T_{\text{c}}=$ 625 kHz. The signaling format of the AU is a $\pi/2$-BPSK. On the other hand, the TU transmits quaternary-phase-shift-keying (QPSK) symbols. The carrier frequency is set to $f_{\text{carrier}}=27$ GHz. Regarding the training protocol, we set $Q_{\text{T}}=16$ and $N_{\text{A, train}}=N_{\text{T, train}}=80$. Unless otherwise specified, the TU coherence time is equal to $N_{\text{coh}}=16.384$ OFDM ksymbols.
The ULA at the BS has $J=4$ antennas. The number of paths of the TU-to-BS link is fixed to $K_{\text{T}}=2$. The Rician factor of the AU channel is $K_{\text{A}}=6$ dB. The direction cosines $\sin(\theta_{\text{A,L}})$ and $\sin(\theta_{\text{A,N}})$ associated with the AoA at the BS are independent random variables uniformly distributed into $[0,1]$. The generic time delay $\tau\in\{\tau_{\text{T},1},\tau_{\text{T},2},\tau_{\text{A,L}},\tau_{\text{A,N}}\}$ of the TU and AU channels is randomly generated according to the one-sided exponentially decreasing delay power spectrum [8], i.e.,
$\tau=-\tau_{\text{slope}}\,\text{ln}\left[1-u\left(1-e^{-\Delta_{\text{max}}/\tau_{\text{slope}}}\right)\right]$ (101)
where $\Delta_{\text{max}}=3\,T_{\text{c}}$, the slope-time is $\tau_{\text{slop}}=2\,T_{\text{c}}$, and $u$ is a random variable uniformly distributed into $[0,1]$. We assume that the constant radial speed of the AU is $v=6$ m/s, $f_{\text{max}}=720$ Hz, and the AoD $\vartheta_{\text{A,L}}$ and $\vartheta_{\text{A,N}}$ are independent random variables uniformly distributed into $[0,\pi]$. The power relationship between the AU and TU at the BS is measured by the aerial-to-terrestrial ratio (ATR) defined as
$\text{ATR}\triangleq\frac{\sigma_{\text{A,L}}^{2}+\sigma_{\text{A,N}}^{2}}{\sum_{k=1}^{K_{\text{T}}}\sigma_{g_{\text{T},k}}^{2}}$ (102)
whereas the signal-to-noise ratio is given by
$\text{SNR}\triangleq\frac{\sigma_{\text{A,L}}^{2}+\sigma_{\text{A,N}}^{2}}{\sigma^{2}_{w}}\>.$ (103)
In each Monte Carlo run, a new set of symbols, noise and channel parameters for both the AU and TU are randomly generated. Regarding the estimation of the channel parameters of the AU (see Section 3), we calculate the arithmetic mean of the MSEs of the Doppler shifts $f_{\text{A,L}}$ and $f_{\text{A,N}}$ (normalized by $f_{\text{max}}^{2}$), the time delays $\tau_{\text{A,L}}$ and $\tau_{\text{A,N}}$ (normalized by $\Delta^{2}_{\text{max}}$), the channel gains $g_{\text{A,L}}$ and $g_{\text{A,N}}$ (normalized by $\sigma^{2}_{\text{A,L}}$), and the direction cosines $\sin(\theta_{\text{A,L}})$ and $\sin(\theta_{\text{A,N}})$. With reference to the estimation of the TU channel (see Section 4), we report the estimator output (97) (normalized by $\sum_{k=1}^{K_{\text{T}}}\sigma_{g_{\text{T},k}}^{2}$).
Figure 2: MSE of the AU Doppler shifts versus the length $N_{0}$ of the observation interval, ($\text{ATR}=0$ dB and $\text{SNR}=14$ dB).
Figure 3: MSE of the AU Doppler shifts versus SNR and ATR values.
Figure 4: MSE of the AU time delays versus SNR and ATR values.
Figure 5: MSE of the AU channel gains versus SNR and ATR values.
Figure 6: MSE of the AU directive cosines versus SNR and ATR values.
Figure 7: MSE of TU channel versus SNR and ATR values (dotted line for the LS estimator and solid line for the BWLU estimator).
Figs. 26 depict the estimation performance of the proposed finite-sample estimators outlined in Section 3 for the AU channel parameters (i.e., Doppler shifts, time delays, channel gains, and AoAs). Generally, it is apparent that such estimators exhibit satisfactory MSE performance for all the considered SNR and ATR values.
Fig. 2 reports the MSE of the Doppler shifts as a function of the length $N_{0}$ of the observation interval (in OFDM symbols). As expected, the estimation accuracy rapidly improves as $N_{0}$ increases, thus demonstrating that the proposed blind estimator of the Doppler shifts is asymptotically consistent. Remarkably, Fig. 3 shows that the MSE of the Doppler shifts is independent of both the SNR and ATR, achieving the value $-62.57$ dB approximately. Such an invariance with respect to both the SNR and ATR comes from the fact that the blind estimator proposed in Subsection III-A-1 relies on the second-order conjugate cyclic statistics of the received data, for which the contribution of the circular wide-sense stationary TU signal and noise is negligible for sufficiently large values of $N_{0}$.
We remember that, as the Doppler shifts, the time delays are blindly estimated by using the conjugate second-order statistics of the ACS random process (17). In this case, a weak dependence of the MSE on the SNR and ATR values can be observed in Fig. 4, which is due to the fact that the estimates of the time delays in Subsection III-A-2 also feel the aftereffects of the estimation error of the cycle frequency acquired previously. However, the variability of the MSE of the time delays is confined in a range of about $1$ dB as a function of the SNR and the MSE curves translate down of nearly $1$ dB further to an ATR reduction of $3$ dB.
The MSE of the channel amplitudes in Fig. 5 also exhibits a weak dependence on the SNR and ATR values. However, the additional exploitation of the pilot symbols allows to achieve MSE values smaller than $-10$ dB for $\text{SNR}>4$ dB, even when the ATR is as low as $-3$ dB. In contrast, the MSE of the AoAs in Fig. 6 is actually independent of the ATR, thus further corroborating the robustness of the proposed pilot-based estimators developed in Subsection III-B against the interference generated by the TU.
Fig. 7 depicts the MSE performance of the TU channel estimator proposed in Section 4. Besides the MSE of the BWLU estimator (98) (solid line), we also report the MSE of the simpler LS estimator (87) (dotted line). As predicted by our analysis, the BWLU estimator largely outperforms the LS one since it additionally exploits the CSI knowledge of the AU. This demonstrates the importance of firstly estimate the channel parameters of the AU and, then, using the obtained estimates to acquire the CSI of the TU.
6 Conclusions
In this paper, we have considered the scenario where a flying AU is paired with a static TU through power-domain uplink NOMA. Despite many existing works, we have investigated the performance of finite-sample channel estimation procedures. We have shown that blind and pilot-aided techniques can be suitably combined to accurately estimate the AU and TU channel that are characterized by different second-order wide-sense properties in both the ACS (or Doppler) and modulation domains. Specifically, the AU mobility induces ACS features in the signal received from sky due to Doppler effects that are not present in the signal received from ground, provided that the TU is static. Moreover, the AU can employ a noncircular modulation format, as opposed to the TU that transmits circular symbols. We have demonstrated that Doppler shifts and time delays of the AU can be estimated blindly with high accuracy that is basically independent of the TU and noise powers. The remaining AU parameters, i.e., channel gains and AoAs, can be instead estimated through a pilot-based LS approach that exploit the previously acquired knowledge of the Doppler shifts and time delays of the AU. By equipping the BS with the CSI knowledge of the AU, the overall channel of the TU can be subsequently acquired through a high-precision pilot-based BWLU estimator.
In BPSK modulation, the symbols $\{b_{\text{A}}[n]\}$ are modeled as a sequence of i.i.d. real-valued random variables, each assuming equiprobable values in {±$\{\pm 1\}$. A variation of BPSK is $\pi/2$-BPSK modulation scheme [40, 41, 42], wherein two sets of BPSK constellations are shifted by $\pi/2$, i.e.,
$s_{\text{A}}[n]=e^{j\frac{\pi}{4}}\,e^{j\frac{\pi}{2}[n]_{2}}\,b_{\text{A}}[n]\>.$ (104)
It can be readily verified that ${\mathbb{E}}(|s_{\text{A}}[n]|^{2})=1$ and ${\mathbb{E}}(s_{\text{A}}^{2}[n])=j\,(-1)^{n}$. Consequently, under the standard assumption that the number of subcarriers $M$ is even, one has ${\mathbb{E}}({\mathbf{s}}_{\text{A}}[n]\,{\mathbf{s}}_{\text{A}}^{\text{H}}[n])={\mathbf{I}}_{M}$, whereas ${\mathbb{E}}({\mathbf{s}}_{\text{A}}[n]\,{\mathbf{s}}_{\text{A}}^{\text{T}}[n])={\mathbf{\Delta}}$, with ${\mathbf{\Delta}}\triangleq j\,{\text{diag}}(1,-1,1,\ldots,-1)\in\mathbb{C}^{M\times M}$.